Find the pH during the titration of 20.00 mL of 0.1000 M butanoic acid, CH CH CH COOH (K, 1.54 x 10), with 0.1000 M NaOH solution after the following additions of titrant. (a) 11.00 mL: pH- (b) 20.40 mL: (c) 29.00 mL: pH =
Answer #1

a)when 11.0 mL of NaOH is added

Given:

M(C3H7COOH) = 0.1 M

V(C3H7COOH) = 20 mL

M(NaOH) = 0.1 M

V(NaOH) = 11 mL

mol(C3H7COOH) = M(C3H7COOH) * V(C3H7COOH)

mol(C3H7COOH) = 0.1 M * 20 mL = 2 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.1 M * 11 mL = 1.1 mmol

We have:

mol(C3H7COOH) = 2 mmol

mol(NaOH) = 1.1 mmol

1.1 mmol of both will react

excess C3H7COOH remaining = 0.9 mmol

Volume of Solution = 20 + 11 = 31 mL

[C3H7COOH] = 0.9 mmol/31 mL = 0.029M

[C3H7COO-] = 1.1/31 = 0.0355M

They form acidic buffer

acid is C3H7COOH

conjugate base is C3H7COO-

Ka = 1.54*10^-5

pKa = - log (Ka)

= - log(1.54*10^-5)

= 4.812

use:

pH = pKa + log {[conjugate base]/[acid]}

= 4.812+ log {3.548*10^-2/2.903*10^-2}

= 4.9

Answer: 4.90

b)

when 20.4 mL of NaOH is added

Given:

M(C3H7COOH) = 0.1 M

V(C3H7COOH) = 20 mL

M(NaOH) = 0.1 M

V(NaOH) = 20.4 mL

mol(C3H7COOH) = M(C3H7COOH) * V(C3H7COOH)

mol(C3H7COOH) = 0.1 M * 20 mL = 2 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.1 M * 20.4 mL = 2.04 mmol

We have:

mol(C3H7COOH) = 2 mmol

mol(NaOH) = 2.04 mmol

2 mmol of both will react

excess NaOH remaining = 0.04 mmol

Volume of Solution = 20 + 20.4 = 40.4 mL

[OH-] = 0.04 mmol/40.4 mL = 0.001 M

use:

pOH = -log [OH-]

= -log (9.901*10^-4)

= 3.0043

use:

PH = 14 - pOH

= 14 - 3.0043

= 10.9957

Answer: 11.00

c)

when 29.0 mL of NaOH is added

Given:

M(C3H7COOH) = 0.1 M

V(C3H7COOH) = 20 mL

M(NaOH) = 0.1 M

V(NaOH) = 29 mL

mol(C3H7COOH) = M(C3H7COOH) * V(C3H7COOH)

mol(C3H7COOH) = 0.1 M * 20 mL = 2 mmol

mol(NaOH) = M(NaOH) * V(NaOH)

mol(NaOH) = 0.1 M * 29 mL = 2.9 mmol

We have:

mol(C3H7COOH) = 2 mmol

mol(NaOH) = 2.9 mmol

2 mmol of both will react

excess NaOH remaining = 0.9 mmol

Volume of Solution = 20 + 29 = 49 mL

[OH-] = 0.9 mmol/49 mL = 0.0184 M

use:

pOH = -log [OH-]

= -log (1.837*10^-2)

= 1.736

use:

PH = 14 - pOH

= 14 - 1.736

= 12.264

Answer: 12.26